OpenStudy (anonymous):

Will give out medals. Ethan is 40 years old. His son is 12 years old. How many years ago was Ethan 5 times older than his son?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

I get 5 ..

OpenStudy (anonymous):

Let the years ago was Ethan 5 times older than his son be x According to question, 40 - x = 5(12 - x) 40 - x = 60 - 5x x = 5

OpenStudy (anonymous):

Let Son's age be x then father has to be 5x, now, father is 40 and son is 12. thus the difference of their ages is 28 years, then 5x-x = 28, or x=7 , Therefore, when the son was 7, his father was 5(7)=35 , which is 5 years back, hence the answer is 5